An enemy spaceship is moving in pursuit of your st

just do a, cant reward more than 300 points, i will make the other part seperate so each will be 300
Answer #1

the speed of the missile as seen from A=(0.7*c+0.4*c)/sqrt(1+(0.7*c*0.4*c/c^2))=0.9727*c

b)time taken by the missile=8*10^6*1000 /(0.9727*3*10^8)=27.415 sec

time taken by the laser=8*10^6*1000 /(3*10^8)=26.667 sec